### v1

# A remark on $\Lambda^{2}$-enlargeable manifolds

Abstract.
In this note, we consider the case that the condition “constant near infinity” in the definition of $\Lambda^{2}$-enlargeable manifold replaced by the condition “locally constant near infinity” and prove that $\Lambda^{2}$-enlargeable manifold in the current sense also can not carry a complete Riemannian metric of positive scalar curvature. As a consequence, we give another proof of Wang-Zhang’s theorem on the generalized Geroch conjecture for complete spin manifolds.
1. Introduction
$\Lambda^{2}$-enlargeable manifold was introduced by Gromov-Lawson ([2]). A famous theorem of Gromov-Lawson ([2]) states that $\Lambda^{2}$-enlargeable manifold can not carry a complete Riemannian metric of positive scalar curvature.
Let $W$ be a closed $\Lambda^{2}$-enlargeable manifold and $M$ be noncompact connected spin manifold without boundary with ${\rm dim}M={\rm dim}W$, Wang and Zhang ([8]) proved that the connected sum M#WM\#W can not carry a complete Riemannian metric of positive scalar curvature using the result in [9]. If $M$ is a closed spin manifold, then M#WM\#W is a closed $\Lambda^{2}$-enlargeable manifold ([2]). So it is natural to ask that whether M#WM\#W is a $\Lambda^{2}$-enlargeable manifold for the case that $M$ is a noncompact spin manifold. For noncompact $M$, from the constructions in [7, 8] (which goes back to [2]), one finds that the condition that the maps from the covering manifolds to the standard sphere are constant near infinity does not satisfy for M#WM\#W. In fact the maps are locally constant near infinity. In this note, we consider the $\Lambda^{2}$-enlargeable manifold in this case and prove that the $\Lambda^{2}$-enlargeable manifold in the current sense also can not carry a complete Riemannian metric of positive scalar curvature.
Definition 1.1 ([2]).
A $C^{1}$-map $\varphi:X\to Y$ between Riemannian manifolds is said to be $(\epsilon,\Lambda^{2})$-contracting, if for all $x\in X$, the map $\varphi_{*}:\Lambda^{2}(T_{x}X)\to\Lambda^{2}(T_{\varphi(x)}Y)$ satisfies
$|\varphi_{*}(V_{x}\wedge W_{x})|\leq\epsilon|V_{x}\wedge W_{x}|$
for any $V_{x},W_{x}\in T_{x}X$.
In the following definition of $\Lambda^{2}$-enlargeable Riemannian metric, we replace “constant near infinity” in [2, Definition 7.1] by “locally constant near infinity”. In [5, Definition 1.10], Shi also considered this case.
Definition 1.2.
A Riemannian metric on a connected manifold $M$ is called $\Lambda^{2}$-enlargeable if given any $\epsilon>0$, there exists a covering manifold $M_{\epsilon}\to M$ such that $M_{\epsilon}$ is spin and a smooth map $f_{\epsilon}:M_{\epsilon}\to S^{{\rm dim}M}(1)$ which is $(\epsilon,\Lambda^{2})$-contracting with respect to the lifted metric, locally constant near infinity and of non-zero degree. A connected (not necessarily compact) manifold is said to be $\Lambda^{2}$-enlargeable if any Riemannian metric (not necessarily complete) on $M$ is $\Lambda^{2}$-enlargeable.
Theorem 1.3.
A manifold $M$ without boundary which is $\Lambda^{2}$-enlargeable in the sense of Definition 1.2, can not carry a complete Riemannian metric of positive scalar curvature.
A proof of Theorem 1.3 will be given in Section 2.
Lemma 1.4.
Let $W$ be a closed $\Lambda^{2}$-enlargeable manifold and $M$ be a noncompact connected spin manifold without boundary with ${\rm dim}M={\rm dim}W$, then M#WM\#W is a $\Lambda^{2}$-enlargeable manifold in the sense of Definition 1.2.
Proof.
Let $g^{T(M\#W)}$ be any Riemannian metric on T(M#W)T(M\#W). We fix a point $p\in W$. For any $r\geq 0$, let $B^{W}_{p}(r)=\{y\in W:d(p,y)\leq r\}$. Let $b_{0}>0$ be a fixed sufficiently small number. Then the connected sum M#WM\#W can be constructed so that the hypersurface $\partial B^{W}_{p}(b_{0})$, which is the boundary of $B^{W}_{p}(b_{0})$, cuts M#WM\#W into two parts: the part $W\setminus B^{W}_{p}(b_{0})$ and the rest part coming from $M$ (by attaching the boundary of a ball in $M$ to $\partial B^{W}_{p}(b_{0})$).
From the metric $g^{T(M\#W)}$, by restriction we get a metric $g^{T(W\setminus B^{W}_{p}(b_{0}))}$ on $T(W\setminus B^{W}_{p}(b_{0}))$, and we extend it to a smooth metric $g^{TW}$ on $TW$. Then $g^{TW}$ is a $\Lambda^{2}$-enlargeable metric. For any $\epsilon>0$, let $\pi:\widehat{W}_{\epsilon}\to W$ be a covering manifold verifying [2, Definition 7.1], carrying lifted geometric data from that of $W$. Especially there exists a smooth map $f:\widehat{W}_{\epsilon}\to S^{{\rm dim}M}(1)$ and a compact subset $K_{\epsilon}\subset\widehat{W}_{\epsilon}$ such that $f$ is constant on $\widehat{W}_{\epsilon}\setminus K_{\epsilon}$. As in [7, 8], the connected sum M#WM\#W lifts naturally to $\widehat{W}_{\epsilon}$ and we denote the resulting manifold by $\widehat{M}\#\widehat{W}_{\epsilon}$. We lift the metric $g^{T(M\#W)}$ to $\widehat{M}\#\widehat{W}_{\epsilon}$. We extend the map $f$ to $\widehat{M}\#\widehat{W}_{\epsilon}$ by seeting that $f|_{\widehat{M}\#(\widehat{W}_{\epsilon}\setminus K_{\epsilon})}$ is constant. Then by the construction in [8], we have a map $\widehat{f}:\widehat{M}\#\widehat{W}_{\epsilon}\to S^{{\rm dim}M}(1)$ which is $(c\epsilon,\Lambda^{2})$-contracting with respect to the lifted metric for some constant $c>0$, locally constant near infinity and of non-zero degree. Then the metric $g^{T(M\#W)}$ is a $\Lambda^{2}$-enlargeable metric in the sense of Definition 1.2 and by definition M#WM\#W is a $\Lambda^{2}$-enlargeable manifold in the sense of Definition 1.2.
By Theorem 1.3 and Lemma 1.4, we get [8, Theorem 1.1], which states that M#WM\#W can not carry a complete Riemannian metric of positive scalar curvature.
2. A proof of Theorem 1.3
Let $g^{TM}$ be a complete Riemannian metric on $TM$ and $k^{TM}$ be the associated scalar curvature. We argue by contradiction. Assume that
$k^{TM}>0\ \ {\rm over}\ \ M.$
Following the proof of [2, Theorem 6.12], we consider another metric on $TM$ defined by $k^{TM}g^{TM}$. By definition, for the metric $k^{TM}g^{TM}$ and any $\epsilon>0$, there exists a covering
$\pi_{\epsilon}:M_{\epsilon}\to M$
such that $M_{\epsilon}$ is spin and a smooth map
$f_{\epsilon}:M_{\epsilon}\to S^{{\rm dim}M}(1)$
which is $(\epsilon,\Lambda^{2})$-contracting for the lifted metric of $k^{TM}g^{TM}$, locally constant outside a compact subset $K_{\epsilon}$ and of non-zero degree.
Let $g^{TM_{\epsilon}}=\pi^{*}_{\epsilon}g^{TM}$ be the lifted metric of $g^{TM}$ and $k^{TM_{\epsilon}}=\pi^{*}(k^{TM})$. Set ${\rm dim}M=n$.
For $M_{\epsilon}$, $g^{TM_{\epsilon}}$ and $f_{\epsilon}:M_{\epsilon}\to S^{n}(1)$, we do the same constructions as in [4, Section 3.1] and we use the same notations except that we add $\epsilon$ to denote the manifold $M_{\epsilon}$. Then we have a compact manifold $M_{H_{\epsilon,3m}}$ with boundary $H_{\epsilon,3m}$ and a closed manifold $\widehat{M}_{H_{\epsilon,3m}}$ with the metric $g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}$. Let $\nabla_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}$ be the associated Levi-Civita connection. The map $f_{\epsilon}|_{M_{H_{\epsilon,3m}}}$ also extends to a map $f_{\epsilon,l}:\widehat{M}_{H_{\epsilon,3m}}\to S^{n}(1)$.
We first assume that $n$ is even.
Let $S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})=S_{\beta,+}(T\widehat{M}_{H_{\epsilon,3m}})\oplus S_{\beta,-}(T\widehat{M}_{H_{\epsilon,3m}})$ be the $\mathbb{Z}_{2}$-graded Hermitian vector bundle of spinors associated to $(T\widehat{M}_{H_{\epsilon,3m}},g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}})$, carrying the canonical induced Hermitian connection $\nabla^{S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})}=\nabla^{S_{\beta,+}(T\widehat{M}_{H_{\epsilon,3m}})}\oplus\nabla^{S_{\beta,-}(T\widehat{M}_{H_{\epsilon,3m}})}$.
Let $S(TS^{n}(1))=S_{+}(TS^{n}(1))\oplus S_{-}(TS^{n}(1))$ be the spinor bundle of $S^{n}(1)$. Let $V:S(TS^{n}(1))\to S(TS^{n}(1))$ be the map defined in [6, Section 2.5]. Then there exists $\delta>0$ such that
(2.1) $(f^{*}_{\epsilon,l}V)^{2}\geq\delta\ \ {\rm on}\ \ \widehat{M}_{H_{\epsilon,3m}}\setminus{\rm Supp}({\rm d}f_{\epsilon}).$
Let
(2.2) $\left(E_{3m,\pm},g^{E_{3m,\pm}},\nabla^{E_{3m,\pm}}\right)=f_{\epsilon,l}^{*}\left(S_{\pm}(TS^{n}(1)),g^{S_{\pm}(TS^{n}(1))},\nabla^{S_{\pm}(TS^{n}(1))}\right)$
be the induced Hermitian vector bundle with the Hermitian connection on $\widehat{M}_{H_{\epsilon,3m}}$. Then $E_{3m}=E_{3m,+}\oplus E_{3m,-}$ is a ${\mathbb{Z}}_{2}$-graded Hermitian vector bundle over $\widehat{M}_{H_{\epsilon,3m}}$.
Let $\nabla^{S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})\widehat{\otimes}E_{3m}}$ be the connection on $S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})\widehat{\otimes}E_{3m}$ induced by $\nabla^{S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})}$ and $\nabla^{E_{3m,\pm}}$.
Let $D^{E_{3m}}_{\beta}$ acting on $S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})\widehat{\otimes}E_{3m}$ be the twisted Dirac operators defined by
(2.3) $D^{E_{3m}}_{\beta}=\sum_{i=1}^{n}c_{\beta}(h_{i})\nabla_{h_{i}}^{S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})\widehat{\otimes}E_{3m}},$
where $\{h_{1},\cdots,h_{n}\}$ is a local oriented orthonormal basis of $(T\widehat{M}_{H_{\epsilon,3m}},g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}})$, and $c_{\beta}(\cdot)$ means that the Clifford action is with respect to the metric $g^{T\widehat{M}_{H_{\epsilon,3m}}}_{\beta}$.
For $\varepsilon>0$, we introduce the following deformation of $D^{E_{3m}}_{\beta}$ on $\widehat{M}_{H_{\epsilon,3m}}$,
(2.4) $D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}},$
and let
(2.5) $\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)_{+}:\Gamma\left(S_{\beta,+}\left(T\widehat{M}_{H_{\epsilon,3m}}\right){\otimes}E_{3m,+}\oplus S_{\beta,-}\left(T\widehat{M}_{H_{\epsilon,3m}}\right){\otimes}E_{3m,-}\right)\\ \to\Gamma\left(S_{\beta,-}\left(T\widehat{M}_{H_{\epsilon,3m}}\right){\otimes}E_{3m,+}\oplus S_{\beta,+}\left(T\widehat{M}_{H_{\epsilon,3m}}\right){\otimes}E_{3m,-}\right)$
be the natural restriction.
By the Lichnerowicz formula, we have
(2.6) $\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)^{2}=\left(D^{E_{3m}}_{\beta}\right)^{2}+\left[D^{E_{3m}}_{\beta},{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right]+\left({{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)^{2}\\ =-\Delta^{E_{3m},\beta}+{{k^{T\widehat{M}_{H_{\epsilon,3m}}}}\over{4}}+{1\over 2}\sum_{i,j=1}^{n}R^{E_{3m}}(h_{i},h_{j})c_{\beta}(h_{i})c_{\beta}(h_{j})+\left[D^{E_{3m}}_{\beta},{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right]+\left({{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)^{2},$
where $\Delta^{E_{3m},\beta}\geq 0$ is the corresponding Bochner Laplacian, $k^{T\widehat{M}_{H_{\epsilon,3m}}}$ is the scalar curvature of $g_{\beta}^{T\widehat{M}_{H_{\epsilon,3m}}}$ and
$R^{E_{3m}}=\left(\nabla^{E_{3m,+}}\right)^{2}+\left(\nabla^{E_{3m,-}}\right)^{2}.$
On $TM_{H_{\epsilon,3m}}$, by the $(\epsilon,\Lambda^{2})$-contracting property of $f_{\epsilon}$ for the metric $k^{TM_{\epsilon}}g^{TM_{\epsilon}}$, we have
(2.7) $|f_{\epsilon,l,*}(h_{i}\wedge h_{j})|={1\over\beta^{2}}|f_{\epsilon,l,*}(\beta h_{i}\wedge\beta h_{j})|\leq{\epsilon\over{\beta^{2}}}|\beta h_{i}\wedge\beta h_{j}|_{k^{TM_{\epsilon}}g^{TM_{\epsilon}}}={{\epsilon k^{TM_{\epsilon}}}\over{\beta^{2}}},\ \ i\neq j.$
Let $\nabla^{S(TS^{n}(1))}$ be the canonical connection on the spinor bundle of $S^{n}(1)$. Let $R^{S(TS^{n}(1))}$ be the curvature tensor of the connection. Set
(2.8) $C_{1}=\sup_{p\in S^{n}(1)}\Big|R_{p}^{S\left(TS^{n}(1)\right)}\Big|.$
For $x\in{\rm Supp}({\rm d}f_{\epsilon})$ and $s\in\Gamma(\widehat{M}_{H_{\epsilon,3m}},S_{\beta}(T\widehat{M}_{H_{\epsilon,3m}})\widehat{\otimes}E_{3m})$, by (2.7), we have
(2.9) $\left|\left({1\over 2}\sum_{i,j=1}^{n}R^{E_{3m,r}}(h_{i},h_{j})c_{\beta}(h_{i})c_{\beta}(h_{j})s,s\right)(x)\right|\\ =\left|\left({1\over 2}\sum_{i,j}^{n}f^{*}_{\epsilon,l,r}(R^{S}(f_{\epsilon,l,r,*}(h_{i}\wedge h_{j})))c_{\beta}(h_{i})c_{\beta}(h_{j})s,s\right)(x)\right|\leq{{\epsilon k^{TM_{\epsilon}}}\over{2\beta^{2}}}n(n-1)C_{1}|s|^{2}(x),$
where $R^{S}$ is the shorthand for $R^{S(TS^{n}(1))}$.
Now, we choose
(2.10) $\epsilon={1\over{4C_{1}(n+1)^{2}}}.$
Then $f_{\epsilon}$ is fixed and ${\rm Supp}({\rm d}f_{\epsilon})$ is a fixed compact set. Hence, we can find $\kappa>0$ such that
(2.11) $k^{TM_{\epsilon}}\geq\kappa\ \ {\rm on}\ \ {\rm Supp}({\rm d}f_{\epsilon}).$
On $\widehat{M}_{H_{\epsilon,3m}}\setminus(({\rm Supp}({\rm d}f_{\epsilon}))\cup(H_{\epsilon,3m}\times[0,1]))$, we have
(2.12) $\left[D^{E_{3m}}_{\beta},{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right]=0,$
and on $H_{\epsilon,3m}\times[0,1]$, we have
(2.13) $\left[D^{E_{3m}}_{\beta},{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right]=O_{m,\epsilon}\left({\varepsilon\over\beta}\right).$
On ${\rm Supp}({\rm d}f_{\epsilon})$}, we have
(2.14) $\left[D^{E_{3m}}_{\beta},{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right]=O_{\epsilon}\left({\varepsilon\over{\beta^{2}}}\right).$
Then using (2.1), (2.6), (2.9) and (2.11)-(2.14), proceeding as in the proof of [6, Lemma 2.1(i)] and [6, Section 2.5], one finds that there exist $c_{0}>0,\varepsilon>0,m>0$ such that when $\beta>0$ is small enough,
(2.15) $\left\|\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)s\right\|_{\beta}\geq{c_{0}\over\beta}\|s\|_{\beta}.$
On the other hand, by the Atiyah-Singer index theorem [1] (cf. [3, Proposition III. 13.8]), as in [6, (2.44)], we have
(2.16) ${\rm ind}\left(\left(D^{E_{3m}}_{\beta}+{{\varepsilon f^{*}_{\epsilon,l}V}\over{\beta}}\right)_{+}\right)=2(-1)^{n\over 2}{\rm deg}(f_{\epsilon})\neq 0.$
Then we get a contradiction.
If $n$ is odd, we consider the composition $f_{\epsilon,l,r}$ of the maps
(2.17) $\widehat{M}_{H_{\epsilon,3m}}\times S^{1}(r)\xrightarrow{f_{\epsilon,l}\times{1\over r}{\rm id}}S^{n}(1)\times S^{1}(1)\xrightarrow{\wedge}S^{n+1}(1),$
where $S^{1}(r)$ is the round circle of radius $r$ with the canonical metric.
Fix $\epsilon$ as (2.10) and set
$\kappa_{0}=\min\left\{k^{TM_{\epsilon}}(x):x\in{\rm Supp}({\rm d}f_{\epsilon})\right\}.$
We choose $r$ large enough such that
(2.18) ${{\sup\left\{|{{\rm d}}f_{\epsilon}|(x),x\in M_{\epsilon}\right\}}\over r}<\epsilon\kappa_{0}.$
Then by combining the method used in the above even dimensional case and [6, Section 3], we can also get a contradiction.
Acknowledgments. The author would like to thank Pengshuai Shi, Changliang Wang, Xiangsheng Wang and Prof. Weiping Zhang for helpful discussions. This work was partially supported by NSFC Grant No. 12425106, NSFC Grant No. 12271266 and Nankai Zhide Foundation.